# Deep Legendre Transform

### v1

@@ Appendix B Technical appendix: additional experiments
Section B.3 presents comprehensive benchmarks comparing our DLT approach against direct supervised learning when analytical conjugates are known. Using identical neural architectures across multiple functions and dimensions $d\in\{2,\ldots,200\}$, we demonstrate that DLT achieves equivalent $L^{2}$ errors without requiring closed-form duals.
Section B.4 evaluates DLT on challenging non-separable convex functions. We test quadratic forms with random matrices, exponential-minus-linear functions, pre-trained ICNNs, and coupled soft-plus functions with quadratic pairwise interactions. These experiments show that DLT maintains accuracy across strongly coupled functions, although training time increases due to the more complex computation tree required for gradient evaluation in non-separable functions.
Section B.5 compares DLT with an alternative algorithm that uses an approximate inverse gradient mapping instead of exact gradients for training. While DLT trains on exact targets using the Fenchel-Young identity with true gradient values, the alternative method must learn an approximate inverse mapping to construct its training targets. We show that although such an approximation is feasible, it fundamentally depends on the quality of the inverse gradient approximation and provides weaker theoretical guarantees.
Section B.6 describes DLT applications to Hamilton-Jacobi PDEs via the Hopf formula. We introduce Time-Parameterized DLT that approximates $(g^{*}+tH)^{*}$ and compare against Deep Galerkin Methods, showing superior accuracy.
@@ B.1 Grid-based methods: definition-based method and Linear Time Legendre Transform
B.1.1 The definition based method
This method implements the Legendre-Fenchel transform directly according to its fundamental mathematical definition. For a function $f(x)$ in $d$ dimensions, the transform is defined as:
@@ B.1.2 The nested formulation
The nested method exploits the mathematical structure of the Legendre-Fenchel transform by reformulating it as a sequence of nested maximization problems. For a function $f(x)$ in $d$ dimensions, the transform can be rewritten in nested form:
$f^{*}(s_{1},s_{2},\dots,s_{d})=\max_{x_{1}}\left\{s_{1}x_{1}+\max_{x_{2}}\left\{s_{2}x_{2}+\dots+\max_{x_{d}}\left\{s_{d}x_{d}-f(x_{1},x_{2},\dots,x_{d})\right\}\dots\right\}\right\}$
@@ B.1.3 Algorithm structure
The LLT algorithm by Lucet1997 efficiently computes $f^{*}(s)$ in $\mathcal{O}(n+m)$ time, where $n$ is the number of data points and $m$ is the number of slopes:
@@
    Convex hull step: Compute the convex hull of the discrete data points $(x_{i},u(x_{i}))$.
@@ B.1.4 Comparative characteristics
    Computational complexity: $\mathcal{O}(dN^{d+1})$ for the nested approach, compared to $\mathcal{O}(N^{2d})$ for the definition based method. Each stage requires $\mathcal{O}(N^{d+1})$ operations, with $d$ total stages.
@@ Results.
Each benchmark table reports performance and accuracy metrics for various input dimensions $d$, using uniform grids with $n=m=10$ points per dimension [^11] How we choose the dual sampling grid. Let $\mathcal{G}_{x}=\{x^{(\alpha)}\}$ be the Cartesian primal grid. For a proper, convex, coercive function $f$ every optimal slope is a sub-gradient of $f$ at some sampled node, i.e. $\partial f(\mathcal{G}_{x})=\bigcup_{\alpha}\partial f(x^{(\alpha)})$ contains all slopes that can ever maximise $\langle s,x\rangle-f(x)$. We therefore compute (analytically when possible, otherwise once on the discrete grid) coordinate-wise bounds $\underline{s}_{i}\leq\partial_{x_{i}}f(x^{(\alpha)})\leq\overline{s}_{i}$ for all $\alpha$ and set the dual grid to the Cartesian product $\mathcal{G}_{s}=\{\underline{s}_{i}+k\,h_{i}\}_{k=0}^{M-1},\;h_{i}=(\overline{s}_{i}-\underline{s}_{i})/(M-1).$ If $\partial f(\mathcal{G}_{x})\subset\mathcal{G}_{s}$ the finite-slope-grid theorem—a $d$-D extension of Lucet’s Prop. 2 Lucet1997; see also HiriartUrruty1980—guarantees $f^{*}_{\mathcal{G}_{x}}(s)=\max_{x\in\mathcal{G}_{x}}\bigl\langle s,x\bigr\rangle-f(x)=f^{*}(s)\qquad\text{for every }s\in\mathcal{G}_{s},$ so further enlarging the slope range cannot change the transform at any sampled point—only the interpolation error between nodes depends on how fine $m$ is chosen. Example: for $f(x)=-\sum_{i=1}^{d}\log x_{i}$ on $x_{i}\in[0.1,5]$ we have $\partial_{x_{i}}f\in[-10,-0.2]$; taking a uniform dual grid over [10,0.2][-10,-0.2] already covers all sub-gradients and yields the exact conjugate values on that grid..
The column $t_{\text{Lucet}}$ (s) reports the runtime of the Nested Lucet method in seconds, while $t_{\text{Def}}$ (s) gives the runtime of the brute-force definition based method (evaluated only for $d\leq 5$ due to computational constraints). The columns MBL and MBD indicate the peak memory usage in megabytes (MB) during execution for the Lucet and Def methods, respectively; $0.0$ reads that less than $0.05$ MB was used. The column max|LD||L{-}D| measures the maximum absolute difference between the Lucet and Def outputs across all grid points. Lastly, max err captures the interpolation error of the Lucet method relative to a reference solution obtained via uniform sampling, and RMSE reports the root mean squared error over the same validation sample. Some columns are indicated as – since direct computations are too expensive to run.
@@
Table B1: Function properties, grid domains ($10^{d}$ total points), and Legendre transforms.
@@
Quad-Linear $f(x)=\frac{\sum x_{i}^{2}+1}{\sum x_{i}+1}$ $0,3]^{d}$ No simple analytical form $\nabla f([0,3]^{d})$
Table B2: Quadratic function benchmark: Lucet vs. Definition based method comparison across dimensions.
@@
Table B3: Benchmark results for the negative log function. Lucet and Direct methods compared across dimensions. MB: active memory (MB); max|LD||L{-}D|: maximum discrepancy between methods.
$d$ $t_{\text{Lucet}}$ (s) $t_{\text{Direct}}$ (s) MBL MBD max|LD||L{-}D| max err RMSE
@@
Table B4: Estimated runtime comparison: Direct vs. Lucet methods with $N=10$ points per dimension. The Direct method exhibits $O(N^{2d})$ complexity while Lucet achieves $O(dN^{d+1})$.
$d$ Points/dim Direct Method (est.) Lucet Method (est.) Speedup
@@
B.2 Softmax-Based approximation for Legendre-Fenchel Transforms
The softmax-based approach, as proposed by Peyré, see Pey and BlondelRouletGoogle24, offers an alternative method for approximating Legendre-Fenchel transforms through a convolution-based formulation. For a convex function $f$ defined on a bounded domain $C\subset\mathbb{R}^{d}$, the Legendre-Fenchel transform can be approximated using:
$f^{*}_{\epsilon}(s)=\epsilon\log\left(\int_{C}\exp\left(\frac{\langle x,s\rangle-f(x)}{\epsilon}\right)\,dx\right)$
where $\epsilon>0$ is a smoothing parameter that controls the accuracy of the approximation. As $\epsilon\to 0$, the softmax approximation $f^{*}_{\epsilon}$ converges to the exact transform $f^{*}$.
@@ B.2 Softmax-Based approximation for Legendre-Fenchel Transforms
$f^{*}_{\epsilon}(s)=Q^{-1}_{\epsilon}\left(\frac{1}{Q_{\epsilon}(f)}*G_{\epsilon}\right)(s)$
where: $G_{\epsilon}(x)=\exp(-\frac{1}{2\epsilon}\|x\|^{2}_{2})$ is a Gaussian kernel, $Q_{\epsilon}(f)(x)=\exp(\frac{1}{2\epsilon}\|x\|^{2}_{2}-\frac{1}{\epsilon}f(x))$, and $Q^{-1}_{\epsilon}(F)(s)=\frac{1}{2}\|s\|^{2}_{2}-\epsilon\log F(s)$.
@@ B.2.2 Empirical results
In our experiments Monte Carlo approach showed a more stable results for performing integration for small values of $\epsilon$, so we focus on it. Our MC implementation uses Sobol sequences to generate low-discrepancy sample points, achieving better coverage of the integration domain compared to purely random sampling. Table B5 presents results for the negative logarithm function using 100,000 samples. The dual grid is the same as in LLT experiments from the previous section, see Table B3.
Experiments confirm $\mathcal{O}(\epsilon)$ convergence to the exact Legendre transform as $\epsilon\to 0$. Peyré’s approach is suitable when an estimate of the transform is required at only one point $y$ in the dual space, but it is still too expensive to iterate over dual grids/samples, especially in high-dimensional setups: $d>4$.
Table B5: Benchmark for negative log function using quasi-Monte Carlo (Sobol) with 100,000 samples across dimensions and noise levels $\epsilon$.
$d$ $\epsilon$ Sample Size Time (s) MB Max Err RMSE
1 0.001 MC(100k) 1.690 6.43 1.98e-01 4.43e-02
0.010 MC(100k) 1.714 6.13 2.17e-01 5.33e-02
0.100 MC(100k) 1.711 6.14 2.21e-01 1.30e-01
0.500 MC(100k) 1.793 6.13 6.38e-01 3.55e-01
2 0.001 MC(100k) 3.833 9.26 3.95e-01 6.54e-02
0.010 MC(100k) 3.936 9.26 4.33e-01 8.75e-02
0.100 MC(100k) 3.361 9.26 4.42e-01 2.39e-01
0.500 MC(100k) 3.365 9.29 1.28e++$0$ 5.20e-01
3 0.001 MC(100k) 23.794 13.16 6.03e-01 8.73e-02
0.010 MC(100k) 29.116 12.46 6.49e-01 1.21e-01
0.100 MC(100k) 16.801 12.46 6.64e-01 3.47e-01
0.500 MC(100k) 16.960 12.46 1.91e++$0$ 6.56e-01
4 0.001 MC(100k) 419.791 18.21 8.29e-01 1.60e-01
0.010 MC(100k) 561.630 18.66 8.65e-01 1.93e-01
0.100 MC(100k) 317.030 20.06 8.85e-01 4.53e-01
0.500 MC(100k) 317.895 20.06 2.55e++$0$ 7.80e-01
Table B6: Benchmark for negative entropy function using quasi-Monte Carlo (Sobol) with 100,000 samples across dimensions and noise levels $\epsilon$.
$d$ $\epsilon$ Sample Size Time (s) MB Max Err RMSE
1 0.001 MC(100k) 4.379 11.61 4.42e-01 1.03e-01
0.010 MC(100k) 4.196 6.15 4.73e-01 1.18e-01
0.100 MC(100k) 4.145 6.13 5.83e-01 2.19e-01
0.500 MC(100k) 4.150 6.14 9.02e-01 4.75e-01
2 0.001 MC(100k) 8.207 9.26 8.87e-01 1.57e-01
0.010 MC(100k) 8.206 9.28 9.47e-01 1.87e-01
0.100 MC(100k) 7.933 9.26 1.17e++$0$ 3.75e-01
0.500 MC(100k) 7.993 9.26 1.80e++$0$ 7.04e-01
3 0.001 MC(100k) 30.088 12.47 1.39e++$0$ 2.29e-01
0.010 MC(100k) 33.914 12.50 1.44e++$0$ 2.60e-01
0.100 MC(100k) 22.726 13.92 1.75e++$0$ 5.26e-01
0.500 MC(100k) 22.507 13.22 2.70e++$0$ 8.98e-01
4 0.001 MC(100k) 403.168 18.61 2.01e++$0$ 4.10e-01
0.010 MC(100k) 524.356 17.96 2.00e++$0$ 4.33e-01
0.100 MC(100k) 301.965 20.07 2.51e++$0$ 7.17e-01
0.500 MC(100k) 306.520 17.91 3.57e++$0$ 1.07e++$0$
@@ B.3 DLT: Comparing different NN architectures
B.3.1 Comparison with direct learning when dual is known
A critical validation of our Deep Legendre Transform approach is to demonstrate that it achieves state-of-the-art performance by comparing it to the direct learning of the convex conjugate in cases where the analytic form of $f^{*}$ is already known. For this comparison, we implemented a comprehensive benchmark that compares DLT [^12] DLT is an implicit approach, effectively, defined by (1.5). against explicit direct supervised learning across several convex functions with known convex conjugates in multiple dimensions.
@@ B.3.2 Benchmark setup
  • Implicit (DLT): Our proposed approach that trains a neural network $g_{\theta}$ to satisfy $g_{\theta}(\nabla f(x))\approx\langle x,\nabla f(x)\rangle-f(x)$.
  • Direct: The approach that directly approximates $f^{*}$ by training $h_{\theta}$ so that $h_{\theta}(y)\approx f^{*}(y)$, which is only possible when $f^{*}$ is known in closed form.
It is important to emphasize that the direct approach serves as our "gold standard" since it has direct access to the target function values. However, this advantage is only available in the special cases where $f^{*}$ has a known analytical expression—an assumption that does not hold in many practical applications. Our goal is to demonstrate that DLT performs identically to this idealized case.
@@
  • MLP_ICNN: Feed-forward network with enforced convexity using only positive weights (two hidden layers with 128 units each)
@@
  • ICNN: Input-convex neural network following Amos2017 with positive z-weights and direct x-skip connections (two hidden layers with 128 units each)
@@
  • Quadratic $f(x)=\frac{\|x\|^{2}}{2}$: Since $\nabla f(x)=x$, we sample $x\sim\mathcal{N}(0,I)$ for the implicit approach, which naturally gives $y=x\sim\mathcal{N}(0,I)$ in dual space for the explicit approach.
@@
We tested these three convex functions with known analytic duals across dimensions $d\in\{2,5,10,20\}$:
Quadratic: $f(x)=\frac{\|x\|^{2}}{2},\quad$ $f^{*}(y)=\frac{\|y\|^{2}}{2}$ (B1)
@@
The quadratic function $f(x)=\|x\|^{2}/2$ is particularly elegant for our comparison as it is self-conjugate ($f^{*}=f$) and has gradient $\nabla f(x)=x$. This means that when sampling $x\sim\mathcal{N}(0,I)$ for the implicit approach, we get $y=\nabla f(x)=x\sim\mathcal{N}(0,I)$ for the dual space as well, making the comparison especially clean.
Table B8 presents the $L^{2}$ errors and training times for both approaches. These benchmark results reveal a key finding: our implicit DLT approach performs essentially identically to direct learning when the convex conjugate is known analytically. The $L^{2}$ errors for both methods are statistically identical.
When the dual validation samples are drawn from the training distribution, the Deep Legendre Transform learns the convex conjugate just as accurately (and almost as fast) as direct supervised learning of $f^{*}$ — even though it never sees the true dual values.
These results validate a key advantage of our approach: DLT achieves the same approximation quality as methods that require knowing the analytical form of $f^{*}$, without actually needing this knowledge. This makes DLT applicable to the wide range of problems where $f^{*}$ does not have a known closed form, while still maintaining state-of-the-art performance in cases where the dual is known.
Table B7: Benchmark results comparing implicit DLT against direct learning with known duals

### v2

@@ Appendix B Technical appendix: additional experiments
Section B.3 presents comprehensive benchmarks comparing our DLT approach against direct supervised learning when analytical conjugates are known. Using identical neural architectures across multiple functions and dimensions $d\in\{2,\ldots,200\}$, we demonstrate that DLT achieves equivalent $L^{2}$ errors without requiring closed-form Legendre transforms.
Section B.4 evaluates DLT on challenging non-separable convex functions. We test quadratic forms with random matrices, exponential-minus-linear functions, pretrained ICNNs, and coupled soft-plus functions with quadratic pairwise interactions. These experiments show that DLT maintains accuracy across strongly coupled functions, although training time increases due to the more complex computation tree required for gradient evaluation in non-separable functions.
Section B.5 compares DLT with an alternative algorithm that uses an approximate inverse gradient mapping instead of exact gradients for training. While DLT trains on exact targets using the Fenchel–Young identity with true gradient values, the alternative method must learn an approximate inverse mapping to construct its training targets. We show that although such an approximation is feasible, it fundamentally depends on the quality of the inverse gradient approximation and provides weaker theoretical guarantees.
Section B.6 describes DLT applications to Hamilton–Jacobi PDEs via the Hopf formula. We introduce Time-Parameterized DLT that approximates $(g^{*}+tH)^{*}$ and compare against Deep Galerkin Methods, showing superior accuracy.
@@ B.1 Grid-based methods: definition-based method and Linear Time Legendre Transform
B.1.1 The definition-based method
The definition-based method implements convex conjugation directly according to the Legendre–Fenchel definition (1.1)
@@ B.1.2 The nested formulation
The nested method exploits the mathematical structure of the Legendre–Fenchel transform by reformulating it as a sequence of nested maximization problems. For a function $f(x)$ in $d$ dimensions, the transform can be rewritten in nested form
$f^{*}(s_{1},s_{2},\dots,s_{d})=\max_{x_{1}}\left\{s_{1}x_{1}+\max_{x_{2}}\left\{s_{2}x_{2}+\dots+\max_{x_{d}}\left\{s_{d}x_{d}-f(x_{1},x_{2},\dots,x_{d})\right\}\right\}\right\}$
@@ B.1.3 Algorithm structure
The LLT algorithm by Lucet (1997) efficiently computes $f^{*}(s)$ in $\mathcal{O}(n+m)$ time, where $n$ is the number of data points and $m$ is the number of slopes:
@@
    Convex hull step: Compute the convex hull of the discrete data points $(x_{i},f(x_{i}))$.
@@ B.1.4 Comparative characteristics
    Computational complexity: $\mathcal{O}(dN^{d+1})$ for the nested approach, compared to $\mathcal{O}(N^{2d})$ for the definition-based method. Each stage requires $\mathcal{O}(N^{d+1})$ operations, with $d$ total stages.
@@ Results.
Each benchmark table reports performance and accuracy metrics for various input dimensions $d$, using uniform grids with $n=m=10$ points per dimension [^11] How we choose the dual sampling grid. Let $\mathcal{G}_{x}=\{x^{(\alpha)}\}$ be the Cartesian primal grid. For a proper, convex, coercive function $f$ every optimal slope is a sub-gradient of $f$ at some sampled node, i.e. $\partial f(\mathcal{G}_{x})=\bigcup_{\alpha}\partial f(x^{(\alpha)})$ contains all slopes that can ever maximise $\langle s,x\rangle-f(x)$. We therefore compute (analytically when possible, otherwise once on the discrete grid) coordinate-wise bounds $\underline{s}_{i}\leq\partial_{x_{i}}f(x^{(\alpha)})\leq\overline{s}_{i}$ for all $\alpha$ and set the dual grid to the Cartesian product $\mathcal{G}_{s}=\{\underline{s}_{i}+k\,h_{i}\}_{k=0}^{M-1},\;h_{i}=(\overline{s}_{i}-\underline{s}_{i})/(M-1).$ If $\partial f(\mathcal{G}_{x})\subseteq\mathcal{G}_{s}$ the finite-slope-grid theorem—a $d$-D extension of Lucet’s Prop. 2 Lucet (1997); see also Hiriart-Urruty (1980)—guarantees $f^{*}_{\mathcal{G}_{x}}(s)=\max_{x\in\mathcal{G}_{x}}\bigl\langle s,x\bigr\rangle-f(x)=f^{*}(s)\qquad\text{for every }s\in\mathcal{G}_{s},$ so further enlarging the slope range cannot change the transform at any sampled point—only the interpolation error between nodes depends on how fine $m$ is chosen. Example: for $f(x)=-\sum_{i=1}^{d}\log x_{i}$ on $x_{i}\in[0.1,5]$ we have $\partial_{x_{i}}f\in[-10,-0.2]$; taking a uniform dual grid over [10,0.2][-10,-0.2] already covers all sub-gradients and yields the exact conjugate values on that grid..
The column $t_{\text{Lucet}}$ (s) reports the runtime of the Nested Lucet method in seconds, while $t_{\text{Def}}$ (s) gives the runtime of the brute-force definition-based method (evaluated only for $d\leq 5$ due to computational constraints). The columns MBL and MBD indicate the peak memory usage in megabytes (MB) during execution for the Lucet and Def methods, respectively; $0.0$ reads that less than $0.05$ MB was used. The column max|LD||L{-}D| measures the maximum absolute difference between the Lucet and Def outputs across all grid points. Lastly, max err captures the interpolation error of the Lucet method relative to a reference solution obtained via uniform sampling, and RMSE reports the root mean squared error over the same validation sample. Some columns are indicated as – since direct computations are too expensive to run.
@@
Table B1: Functions, grid domains ($10^{d}$ total points) and Legendre transforms.
@@
Quad-Linear $f(x)=\frac{\sum x_{i}^{2}+1}{\sum x_{i}+1}$ $0,3]^{d}$ No explicit form $\nabla f([0,3]^{d})$
Table B2: Quadratic function benchmark: Lucet vs. definition-based method for different dimensions.
@@
Table B3: Benchmark results for the negative log function. Lucet and definition-based methods compared across dimensions. MB: active memory (MB); max|LD||L{-}D|: maximum discrepancy between methods.
$d$ $t_{\text{Lucet}}$ (s) $t_{\text{Def}}$ (s) MBL MBD max|LD||L{-}D| max err RMSE
@@
Table B4: Estimated runtime comparison: definition-based vs. Lucet methods with $N=10$ points per dimension. The definition-based method has $O(N^{2d})$ complexity while Lucet achieves $O(dN^{d+1})$.
$d$ Points/dim Definition-based method (est.) Lucet method (est.) Speedup
@@
B.2 Softmax-Based approximation of Legendre–Fenchel transforms
The softmax-based approach, as proposed by Peyré, see Peyré (2020) and Blondel and Roulet (2024), offers an alternative method for approximating Legendre–Fenchel transforms through a convolution-based formulation. For a convex function $f$ defined on a bounded domain $C\subseteq\mathbb{R}^{d}$, the Legendre–Fenchel transform can be approximated using:
$f^{*}_{\varepsilon}(s)=\varepsilon\log\left(\int_{C}\exp\left(\frac{\langle x,s\rangle-f(x)}{\varepsilon}\right)\,dx\right)$
where $\varepsilon>0$ is a smoothing parameter that controls the accuracy of the approximation. As $\varepsilon\to 0$, the softmax approximation $f^{*}_{\varepsilon}$ converges to the exact transform $f^{*}$.
@@ B.2 Softmax-Based approximation of Legendre–Fenchel transforms
$f^{*}_{\varepsilon}(s)=Q^{-1}_{\varepsilon}\left(\frac{1}{Q_{\varepsilon}(f)}*G_{\varepsilon}\right)(s)$
where: $G_{\varepsilon}(x)=\exp(-\frac{1}{2\varepsilon}\|x\|^{2}_{2})$ is a Gaussian kernel, $Q_{\varepsilon}(f)(x)=\exp(\frac{1}{2\varepsilon}\|x\|^{2}_{2}-\frac{1}{\varepsilon}f(x))$ and $Q^{-1}_{\varepsilon}(F)(s)=\frac{1}{2}\|s\|^{2}_{2}-\varepsilon\log F(s)$.
@@ B.2.2 Empirical results
In our experiments Monte Carlo approach showed a more stable results for performing integration for small values of $\varepsilon$, so we focus on it. Our MC implementation uses Sobol sequences [^12] We tested several MC variants (basic MC / Sobol-QMC / importance / adaptive); Sobol-QMC was most stable in this setting. to generate low-discrepancy sample points, achieving better coverage of the integration domain compared to purely random sampling. Table B5 presents results for the negative logarithm function using 100,000 samples. The dual grid is the same as in LLT experiments from the previous section, see Table B3.
Experiments confirm $\mathcal{O}(\varepsilon)$ convergence to the exact Legendre transform as $\varepsilon\to 0$. Peyré’s approach is suitable when an estimate of the transform is required at only one point $y$ in the dual space, but it is still too expensive to iterate over dual grids/samples, especially in high-dimensional setups: $d>4$.
Table B5: Results of Peyré for negative log function using quasi-Monte Carlo (Sobol) with 100,000 samples across dimensions and smoothing parameter $\varepsilon$.
$d$ $\varepsilon$ Sample Size Time (s) MB Max Err RMSE
1 0.001 qMC(100k) 1.690 6.43 1.98e-01 4.43e-02
0.010 qMC(100k) 1.714 6.13 2.17e-01 5.33e-02
0.100 qMC(100k) 1.711 6.14 2.21e-01 1.30e-01
0.500 qMC(100k) 1.793 6.13 6.38e-01 3.55e-01
2 0.001 qMC(100k) 3.833 9.26 3.95e-01 6.54e-02
0.010 qMC(100k) 3.936 9.26 4.33e-01 8.75e-02
0.100 qMC(100k) 3.361 9.26 4.42e-01 2.39e-01
0.500 qMC(100k) 3.365 9.29 1.28e++$0$ 5.20e-01
3 0.001 qMC(100k) 23.794 13.16 6.03e-01 8.73e-02
0.010 qMC(100k) 29.116 12.46 6.49e-01 1.21e-01
0.100 qMC(100k) 16.801 12.46 6.64e-01 3.47e-01
0.500 qMC(100k) 16.960 12.46 1.91e++$0$ 6.56e-01
4 0.001 qMC(100k) 419.791 18.21 8.29e-01 1.60e-01
0.010 qMC(100k) 561.630 18.66 8.65e-01 1.93e-01
0.100 qMC(100k) 317.030 20.06 8.85e-01 4.53e-01
0.500 qMC(100k) 317.895 20.06 2.55e++$0$ 7.80e-01
Table B6: Results of Peyré for negative entropy function using quasi-Monte Carlo (Sobol) with 100,000 samples across dimensions and smoothing parameter $\varepsilon$.
$d$ $\varepsilon$ Sample Size Time (s) MB Max Err RMSE
1 0.001 qMC(100k) 4.379 11.61 4.42e-01 1.03e-01
0.010 qMC(100k) 4.196 6.15 4.73e-01 1.18e-01
0.100 qMC(100k) 4.145 6.13 5.83e-01 2.19e-01
0.500 qMC(100k) 4.150 6.14 9.02e-01 4.75e-01
2 0.001 qMC(100k) 8.207 9.26 8.87e-01 1.57e-01
0.010 qMC(100k) 8.206 9.28 9.47e-01 1.87e-01
0.100 qMC(100k) 7.933 9.26 1.17e++$0$ 3.75e-01
0.500 qMC(100k) 7.993 9.26 1.80e++$0$ 7.04e-01
3 0.001 qMC(100k) 30.088 12.47 1.39e++$0$ 2.29e-01
0.010 qMC(100k) 33.914 12.50 1.44e++$0$ 2.60e-01
0.100 qMC(100k) 22.726 13.92 1.75e++$0$ 5.26e-01
0.500 qMC(100k) 22.507 13.22 2.70e++$0$ 8.98e-01
4 0.001 qMC(100k) 403.168 18.61 2.01e++$0$ 4.10e-01
0.010 qMC(100k) 524.356 17.96 2.00e++$0$ 4.33e-01
0.100 qMC(100k) 301.965 20.07 2.51e++$0$ 7.17e-01
0.500 qMC(100k) 306.520 17.91 3.57e++$0$ 1.07e++$0$
@@ B.3 DLT: Comparing different NN architectures
B.3.1 Comparison with direct learning when the Legendre transform is known
A critical validation of DLT is to demonstrate that it achieves state-of-the-art performance by comparing it to the direct learning of the convex conjugate in cases where the analytic form of $f^{*}$ is already known. For this comparison, we implemented a comprehensive benchmark that compares DLT [^13] DLT is an implicit approach, effectively, defined by (1.5). against explicit direct supervised learning across several convex functions with known convex conjugates in multiple dimensions.
@@ B.3.2 Benchmark setup
  • DLT: trains a neural network $g_{\theta}$ to satisfy $g_{\theta}(\nabla f(x))\approx\langle x,\nabla f(x)\rangle-f(x)$.
  • Direct: directly trains $g_{\theta}$ so that $g_{\theta}(y)\approx f^{*}(y)$, which is only possible if $f^{*}$ is known explicitly.
The direct approach serves as our “gold standard" since it has direct access to the target function values. However, this advantage is only available in the special cases where $f^{*}$ has a known analytical expression—an assumption that does not hold in many practical applications. Our goal is to demonstrate that DLT matches the performance of this idealized case.
@@
  • MLP-ICNN: Feed-forward network with enforced convexity using only positive weights (two hidden layers with 128 units each)
@@
  • ICNN: Input-convex neural network of Amos et al. (2017) with positive z-weights and direct x-skip connections (two hidden layers with 128 units each)
@@
  • Quadratic $f(x)=\frac{\|x\|^{2}_{2}}{2}$: Since $\nabla f(x)=x$, we sample $x\sim\mathcal{N}(0,I)$ for the implicit approach, which naturally gives $y=x\sim\mathcal{N}(0,I)$ in the dual space for the explicit approach.
@@
We tested these three convex functions with known analytic conjugates across dimensions $d\in\{2,5,10,20\}$:
Quadratic: $f(x)=\frac{\|x\|^{2}_{2}}{2},\quad$ $f^{*}(y)=\frac{\|y\|^{2}_{2}}{2}$
@@
The quadratic function $f(x)=\|x\|^{2}_{2}/2$ is particularly elegant for our comparison since it is self-conjugate ($f^{*}=f$) and has gradient $\nabla f(x)=x$. This means that when using DLT with samples $x\sim\mathcal{N}(0,I)$, we get $y=\nabla f(x)=x\sim\mathcal{N}(0,I)$ in the dual space as well, making the comparison especially clean.
Table B7 presents $L^{2}$ errors and training times for both approaches. These benchmark results reveal a key finding: DLT matches the performance of direct learning if the convex conjugate is known analytically. The $L^{2}$ errors for both methods are statistically identical.
When the dual validation samples are drawn from the training distribution, DLT learns the convex conjugate just as accurately (and almost as fast) as direct supervised learning of $f^{*}$ — even though it never sees the true dual values.
These results validate a key advantage of our approach: DLT achieves the same approximation accuracy as methods that require knowing the analytical form of $f^{*}$, without actually needing this knowledge. This makes DLT applicable to the wide range of problems where $f^{*}$ does not have a known closed form, while still maintaining state-of-the-art performance in cases where the Legendre transform is known.
Table B7: Comparison of DLT with direct learning for known Legendre transforms